# Quantum State Generation with Structure-Preserving Diffusion Model

### v1

@@ 3.2 Mirror Map for Complex-Valued Density Matrices
The von Neumann entropy (e.g., Bengtsson & Życzkowski, 2017) is an extension of the concept of Shannon entropy to quantum states. For $X\in\mathbb{C}_{n}^{+}$, the negative von Neumann entropy is defined as,
$\phi(X)=\operatorname{Tr}(X\log X).$ (5)
The mirror map and its inverse are computed analytically as
@@ 4 Experiments
More precisely, we will consider three labeled classes of data, corresponding to quantum mixed states of a $4-$qubit system with three different levels of entanglements. We will first conduct unconditional generation, from the distribution of the union of all classes. Then we will perform conditional generation, from the same union distribution but conditioned on the class label; results similar to that of individually generating from each class would be seen. Finally, we will investigate conditional generation beyond the known class labels, i.e., the ability to interpolate/extrapolate across existing classes in the training data.
@@ 4.1 Problem Setup
To demonstrate the efficacy of our method, we consider an example of generating quantum states based on training data consisting of quantum states with three different levels of entanglement between qubits.
@@
$U_{12}=M_{12}\otimes\mathbb{I}_{4\times 4},\quad U_{34}=\mathbb{I}_{4\times 4}\otimes M_{34}$
where $\mathbb{I}_{n\times n}$ is the identity matrix.
@@
Fully entangled state is a quantum state of the system that corresponds to a situation where there is simultaneous entanglement between any pair of qubits. Each matrix $U_{ij}$ corresponds to the entanglement between only two qubits, thus the matrix is created using complex unitary matrices of dimension $4$.
@@
$U_{13}=\operatorname{Perm}_{23}(M_{13}\otimes\mathbb{I}_{4\times 4}),\quad$ $U_{12}=M_{12}\otimes\mathbb{I}_{4\times 4},$
$U_{14}=\operatorname{Perm}_{13}(\mathbb{I}_{4\times 4}\otimes M_{14}),\quad$ $U_{23}=\mathbb{I}_{2\times 2}\otimes M_{23}\otimes\mathbb{I}_{2\times 2},$
$U_{24}=\operatorname{Perm}_{23}(\mathbb{I}_{4\times 4}\otimes M_{24}),\quad$ $U_{34}=\mathbb{I}_{4\times 4}\otimes M_{34},$
where $\operatorname{Perm}_{ij}$ permutes the rows and columns that correspond to qubit $i$ and $j$, $\mathbb{I}_{n\times n}$ is the identity matrix.
4.2 Data Preparation
$X=\frac{1}{\operatorname{Tr}(Y)}Y.$ (10)
@@ 4.2 Data Preparation
[…]
Unconditional Mixture $0.00054$ $0.00094$ 0.020 0.000076 0.0000110
Product $0.00052$ 0.00079 0.011 0.000070 0.0000025
Conditional Pairwisely $0.00064$ 0.00100 0.018 0.000082 0.0000061
Fully $0.00056$ 0.00092 0.026 0.000100 0.0000062
Magnitude reference 0.0854 0.1554 1.7484 1.1236 0.0134
Sampling a probability distribution on $U(n)$, such as the Haar measure, is also an interesting task. One may view it as a sampling problem on a constrained set, but note $U(n)$ is actually a non-convex constrained set, and thus common constrained sampling tools (e.g., projected Langevin (Bubeck et al., 2015), sophisticated MCMC walkers (Lovász & Simonovits, 1993; Lovász, 1999; Kannan & Narayanan, 2009; Chen et al., 2018), mirror Langevin (Zhang et al., 2020; Li et al., 2022)) may not suit well. Fortunately, due to the Lie group structure of $U(n)$, the problem can be solved by considering the dynamics,
where $\{e_{1},\dots,e_{n^{2}}\}\text{ spans }\mathfrak{u}(n)$, $\mathrm{d}W_{1},\dots,\mathrm{d}W_{n^{2}}$ are independent one dimensional Brownian motion. One notable property of the diffusion process described in (11) is that it converges to a unique invariant distribution and the $g$-marginal of its invariant distribution is the Haar measure of $U(n)$ (Kong & Tao, 2024). Therefore, by numerically integrating (11), we can approximately sample uniformly from the unitary group $U(4)$ to generate the quantum entanglement matrices. This numerical integration can be done computational-efficiently with $g$ exactly remaining on $U(4)$ via the operator splitting approach (Tao & Ohsawa, 2020).
4.3 Evaluation
Distributional-wise metric. To evaluate the similarity between high-dim probability distributions, we mainly use Sliced Wasserstein Distance (SWD) (Kolouri et al., 2019), Maximum Sliced Wasserstein Distance (MSWD) (Deshpande et al., 2019), and Maximum Mean Discrepancies (MMD) (Gretton et al., 2012), in addition to the standard Wasserstein distance, as they suffer less from the curse of dimensionality. Within the family of MMDs, we consider Energy-MMD, which is computed with kernel $k(x,y)=-\|x-y\|_{2}$. We also compute the 1-Wasserstein distance based on values of the physically meaningful low-dim observable of entanglement negativity.
4.4 Training
We train on a dataset of $900,000$ samples, distributed equally across all three classes of quantum states. The entire dataset is first transformed to the dual space by applying the mirror map. Then a DDPM is trained on the transformed dataset. We train for $1,000,000$ iterations with AdamW optimizer, using an initial learning rate of $10^{-3}$ and decay rate of $0.995$ every $1000$ AdamW steps. We summarize our neural network architecture in Appendix.A.3. We adopt the same data post-processing and metrics evaluation used in (Chen et al., 2023), and we elaborate them in Appendix.A.4.
4.5 Results
SPDM accurately learns the distribution of quantum states. Table 1 summarizes our quantitative results for both conditional and unconditional generation of three classes of quantum states. Figure 2 and Figure 3 depict the generated samples compared against ground truth. Our method manages to learn accurately the correct distribution of eigenvalues, which is extremely sensitive to all the entries in the matrix. We note that product states, pairwisely entangled states, and fully entangled states considered in this work share an identical distribution of eigenvalues since the spectrum of matrices is invariant under unitary conjugations. However, they have distinct marginal distributions for each entry due to the quantum entanglement. Our method successfully predicts the correct joint distribution of all the entries in the density matrix, which is also confirmed by small values in all distribution-based metrics listed in Table 1.
@@ 4.5 Results
5 Conclusion
We developed the first diffusion-based method for the generative modeling of quantum states. Our method hard codes physical knowledge into the generative models and generates samples that strictly satisfy the structural constraints of quantum states. We demonstrate that the model accurately learns and reproduces the distribution of quantum states represented by high dimensional density matrix, and can even create physically-meaningful quantum states of unseen level of entanglements, via conditional generation. In the future, it will be an interesting task to design neural network architectures for parametrizing the score function that specialize in quantum state generation for even better generation quality.
6 Impact Statement
This paper presents work whose goal is to advance the field of machine learning and quantum sciences. There could be potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Vanilla SGM on Quantum State Dataset
In our research, we demonstrate that conventional score-based generative models (SGMs) fall short in reliably generating high-dimensional quantum datasets. Figure 5 exemplifies that, although these SGMs can precisely mimic the distribution of a 1-qubit quantum state (4 dimensions), they face significant challenges due to the curse of dimensionality when the dimensionality is expanded to a 4-qubit state (256 dimensions), as highlighted in Figure 6. This limitation likely stems from the current neural network parameterization strategies. Identifying an efficient network architecture capable of effectively capturing the essence of quantum states remains a pivotal, yet unresolved, question, which we aim to address in future work.
@@ A.3 Training and neural network
A.4 Evaluation
@@
The 1-Wasserstein Distance suffers from the curse of dimensionality seriously which is pointed out in Appendix.C in (Chen et al., 2023). Therefore, we use various metrics to comprehensively evaluate the distance between high dimensional distributions. For the scalar distribution of negativity, we stick to using 1-Wasserstein distance. In the main results, when evaluating the sample quality of generated quantum states (which is a density of $512$ dimensions), we use the Sliced-Wasserstein Distance (SWD) type and Maximum Mean Discrepancies (MMD) type of metrics as our major criterion. These metrics include Energy Maximum Mean Discrepancy (Energy-MMD), Max-sliced Wasserstein distance (MSWD), and Sliced-Wasserstein Distance (SWD). Here, Energy-MMD is the MMD distance computed with kernel $k(x,y)=-\|x-y\|_{2}$. Our implementation is adapted from Geoloss ($W_{1}$, Energy-MMD) and POT (SWD, MSWD).

### v2

@@ 3.2 Mirror Map for Complex-Valued Density Matrices
The von Neumann entropy [e.g., 3] is an extension of the concept of Shannon entropy to quantum states. For $X\in\mathbb{C}_{n}^{+}$, the negative von Neumann entropy is defined as $\phi(X)=\operatorname{Tr}(X\log X)$. The mirror map and its inverse are computed analytically as
@@ 4 Experiments
More precisely, we will consider three labeled classes of data, corresponding to quantum mixed states of a $4-$qubit system with three different levels of entanglements. We will first conduct unconditional generation, from the distribution of the union of all classes. Then we will perform conditional generation, from the same union distribution but conditioned on the class label; results similar to that of individually generating from each class would be seen. Finally, we will investigate conditional generation beyond the known class labels, i.e., the ability to interpolate/extrapolate across existing classes in the training data. We defer the details of training, data preparation, and evaluation to appendix.A.5,appendix.A.3,appendix.A.4.
@@ 4.1 Problem Setup
To show the efficacy of our method, we consider an example of generating quantum states based on training data consisting of quantum states with three different levels of entanglement between qubits.
@@
$U_{12}=M_{12}\otimes\mathbb{I}_{4\times 4},\quad U_{34}=\mathbb{I}_{4\times 4}\otimes M_{34}.$
@@
Fully entangled state corresponds to a situation where there is simultaneous entanglement between any pair of qubits. Each matrix $U_{ij}$ corresponds to the entanglement between only two qubits, thus the matrix is created using complex unitary matrices of dimension $4$.
@@
$U_{12}=M_{12}\otimes\mathbb{I}_{4\times 4},$ $U_{13}=\operatorname{Perm}_{23}(M_{13}\otimes\mathbb{I}_{4\times 4}),$ $U_{14}=\operatorname{Perm}_{13}(\mathbb{I}_{4\times 4}\otimes M_{14}),$
$U_{23}=\mathbb{I}_{2\times 2}\otimes M_{23}\otimes\mathbb{I}_{2\times 2},$ $U_{24}=\operatorname{Perm}_{23}(\mathbb{I}_{4\times 4}\otimes M_{24}),$ $U_{34}=\mathbb{I}_{4\times 4}\otimes M_{34},$
where $\operatorname{Perm}_{ij}$ permutes rows and columns corresponding to qubit $i$ and $j$, $\mathbb{I}_{n\times n}$ is the identity matrix.
4.2 Results
SPDM accurately learns the distribution of quantum states. Table 1 summarizes our quantitative results for both conditional and unconditional generation of three classes of quantum states. Figure 2 and Figure 3 depict the generated samples against ground truth. Our method manages to learn accurately the correct distribution of eigenvalues, which is extremely sensitive to all matrix entries. Note that product states, pairwisely entangled states, and fully entangled states considered in this work share an identical eigenvalue distribution since matrix spectrum is invariant under unitary conjugations. However, they have distinct marginal distributions for each entry due to the quantum entanglement. Our method successfully predicts the joint distribution of all the entries in the density matrix, which is also confirmed by small values in all distribution-based metrics listed in Table 1.
[…]
Uncond. Mixture $0.00054$ $0.00094$ 0.020 0.000076 0.0000110
Product $0.00052$ 0.00079 0.011 0.000070 0.0000025
Cond. Pairwisely $0.00064$ 0.00100 0.018 0.000082 0.0000061
Fully $0.00056$ 0.00092 0.026 0.000100 0.0000062
Mag. ref. 0.0854 0.1554 1.7484 1.1236 0.0134
@@ 4.2 Results
5 Conclusion, Limitation, and Future Work
We developed the first diffusion-based method for generative modeling quantum states. It hard codes physical knowledge into the learning models and generates samples that strictly satisfy the structural constraints of quantum states, resulting in accurate generation represented by density matrices. It can even create physically-meaningful states of unseen level of entanglements, via conditional generation.
In the future, it will be interesting to design neural network architectures for parametrizing the score function that specialize in quantum state generation. SOTA scalability is already demonstrated, but to further improve, generic score parametrization may be insufficient. The training procedure has also not been the focus, but there is also significant potential for improvement (e.g., [33]). Finally, this work mainly focuses on learning methodology development, and physical applications to real quantum problems will also be left as an important future direction.
@@ A.1 Vanilla SGM on Quantum State Dataset
In our research, we demonstrate that conventional score-based generative models (SGMs) fall short in reliably generating high-dimensional quantum datasets. Figure 5 exemplifies that, although these SGMs can precisely mimic the distribution of a 1-qubit quantum state (4 dimensions), they face significant challenges due to the curse of dimensionality when the dimensionality is expanded to a 4-qubit state (256 dimensions), as highlighted in Figure 6. Note that in Figure 6, the compared distribution is the true data distribution (in blue) with the generated data distribution (in red). While the blue points seem to be a single point dirac distribution, in fact it has a similar structure as the true distribution depicted in Figure 5. The true distribution in Figure 6 looks like a single point due to the drastic failure of Euclidean SGM on this task and an exploded data scale of the generated samples. This limitation likely stems from the current neural network parameterization strategies. Identifying an efficient network architecture capable of effectively capturing the essence of quantum states remains a pivotal, yet unresolved, question, which we aim to address in future work.
@@ A.3 Training and neural network
We train on a dataset of $900,000$ samples, distributed equally across all three classes of quantum states. The entire dataset is first transformed to the dual space by applying the mirror map. Then a DDPM is trained on the transformed dataset. We train for $1,000,000$ iterations with AdamW optimizer, using an initial learning rate of $10^{-3}$ and decay rate of $0.995$ every $1000$ AdamW steps.
@@
All the experiments are performed on one RTX 4090 and one RTX 3090.
A.4 Data Preparation
$X=\frac{1}{\operatorname{Tr}(Y)}Y.$ (9)
Sampling a probability distribution on $U(n)$, such as the Haar measure, is also an interesting task. One may view it as a sampling problem on a constrained set, but note $U(n)$ is actually a non-convex constrained set, and thus common constrained sampling tools (e.g., projected Langevin [5], sophisticated MCMC walkers [43, 42, 32, 10], mirror Langevin [62, 38]) may not suit well. Fortunately, due to the Lie group structure of $U(n)$, the problem can be solved by considering the dynamics,
where $\{e_{1},\dots,e_{n^{2}}\}\text{ spans }\mathfrak{u}(n)$, $\mathrm{d}W_{1},\dots,\mathrm{d}W_{n^{2}}$ are independent one dimensional Brownian motion. One notable property of the diffusion process described in (10) is that it converges to a unique invariant distribution and the $g$-marginal of its invariant distribution is the Haar measure of $U(n)$ [36]. Therefore, by numerically integrating (10), we can approximately sample uniformly from the unitary group $U(4)$ to generate the quantum entanglement matrices. This numerical integration can be done computational-efficiently with $g$ exactly remaining on $U(4)$ via the operator splitting approach [53].
A.5 Evaluation
Distributional-wise metric. To evaluate the similarity between high-dim probability distributions, we mainly use Sliced Wasserstein Distance (SWD) [35], Maximum Sliced Wasserstein Distance (MSWD) [12], and Maximum Mean Discrepancies (MMD) [21], in addition to the standard Wasserstein distance, as they suffer less from the curse of dimensionality. Within the family of MMDs, we consider Energy-MMD, which is computed with kernel $k(x,y)=-\|x-y\|_{2}$. We also compute the 1-Wasserstein distance based on values of the physically meaningful low-dim observable of entanglement negativity.
Remark A.1.
@@ A.5 Evaluation
The 1-Wasserstein Distance suffers from the curse of dimensionality seriously which is pointed out in Appendix.C in [9]. Therefore, we use various metrics to comprehensively evaluate the distance between high dimensional distributions. For the scalar distribution of negativity, we stick to using 1-Wasserstein distance. In the main results, when evaluating the sample quality of generated quantum states (which is a density of $512$ dimensions), we use the Sliced-Wasserstein Distance (SWD) type and Maximum Mean Discrepancies (MMD) type of metrics as our major criterion. These metrics include Energy Maximum Mean Discrepancy (Energy-MMD), Max-sliced Wasserstein distance (MSWD), and Sliced-Wasserstein Distance (SWD). Here, Energy-MMD is the MMD distance computed with kernel $k(x,y)=-\|x-y\|_{2}$. Our implementation is adapted from Geoloss ($W_{1}$, Energy-MMD) and POT (SWD, MSWD).
A.6 Additional Numerical Experiments
In this section, we provide additional plots for experiments on $6-$ qubits. The description of these results can be found in the caption of each figure.
Figure 7: Unconditional generation of product states with $6$ qubits. Each qubit is i.i.d. sampled with eigenvalues bounded between $1$ and $3$. Each figure shows an observable for comparison between unconditionally generated samples and their corresponding ground truth samples. From Left to Right: (1) Real value of matrix entries in the dual space ($\operatorname{Real}(X_{1,1})$ v.s. $\operatorname{Real}(X_{2,2}))$ (2) Real value of matrix entries in the dual space ($\operatorname{Real}(X_{3,3})$ v.s. $\operatorname{Real}(X_{4,4}))$ (3) Real value of matrix entries in the primal space ($\operatorname{Real}(X_{1,1})$ v.s. $\operatorname{Real}(X_{2,2}))$ (4) Real value of matrix entries in the primal space ($\operatorname{Real}(X_{3,3})$ v.s. $\operatorname{Real}(X_{4,4}))$
Figure 8: Unconditional generation of pairwisely-entangled states with $6$ qubits. The product states are generated as tensor products of $6$ single qubit quantum state density matrix. Each qubit is i.i.d. sampled with eigenvalues bounded between $1$ and $3$. The pairwisely entangled states are generated from the product states by imposing entanglement on qubit pairs $\{1,2\},\{3,4\},\{5,6\}$. The entanglement matrices are sampled from the Haar measure of $U(4)$. Each figure shows an observable for comparison between unconditionally generated samples and their corresponding ground truth samples. From Left to Right: (1) Real value of matrix entries in the dual space ($\operatorname{Real}(X_{1,1})$ v.s. $\operatorname{Real}(X_{2,2}))$ (2) Real value of matrix entries in the dual space ($\operatorname{Real}(X_{3,3})$ v.s. $\operatorname{Real}(X_{4,4}))$ (3) Real value of matrix entries in the primal space ($\operatorname{Real}(X_{1,1})$ v.s. $\operatorname{Real}(X_{2,2}))$ (4) Real value of matrix entries in the primal space ($\operatorname{Real}(X_{3,3})$ v.s. $\operatorname{Real}(X_{4,4}))$
Figure 9: Unconditional generation of maximally-entangled quantum states with $4$ qubits. Entanglement matrix $U$ are sampled from Haar measure of $U(16)$ instead of pairwisely from all pairs of qubits. Each figure shows an observable for comparison between unconditionally generated samples and their corresponding ground truth samples. From Left to Right: (1) Real value of matrix entries in the dual space ($\operatorname{Real}(X_{1,1})$ v.s. $\operatorname{Real}(X_{2,2}))$ (2) Real value of matrix entries in the dual space ($\operatorname{Real}(X_{3,3})$ v.s. $\operatorname{Real}(X_{4,4}))$ (3) Real value of matrix entries in the primal space ($\operatorname{Real}(X_{1,1})$ v.s. $\operatorname{Real}(X_{2,2}))$ (4) Real value of matrix entries in the primal space ($\operatorname{Real}(X_{3,3})$ v.s. $\operatorname{Real}(X_{4,4}))$
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: we provided the scope and contribution in the abstract.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: please see the Conclusion section in the main text.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide the full proof and the certain assumptions.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provided most of the details in order to reproduce the experiments in the Appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We planned to release the code upon the acceptance of the paper. Also, we have included most of the details in the Appendix for reproducing the experiments.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provided the hyperparameters optimizer etc. in the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: we have reported the numerical results under various metrics and provided comparable baselines.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the hardware setup in the Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We are aligned with NeurIPS Ethics Guidelines.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: This is a proof of concept paper. Broader Impacts is not applicable.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This is a proof of concept paper. The safeguards is not applicable.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The code is original. For the data, we disclosed the original owners.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: It is not applicable for our paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This is no human subjects included in this paper.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This is not applicable for this paper.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.